There is a collection of nickles, dimes, and quarters with the total value $52.10. There are 56 more dimes than nickles. There are 3 times as many quarters as dimes. How many quarters are in the collection?
Answer by mananth(16946)

nickels x
dimes x+56
quarters3(x+56)

Value 52.1

5x+10(x +56)+25 *3(x+56)=5210

5x+10x+560+75x+4200=5210

90x+4760=5210

90x=5210-4760
90x=450
/90
x=5 nickels
dimes= 61
quarters=183